I cannot figure out how to set up and solve this word problem. PLEASE HELP!!!!
Here is goes:
"One half of a number is 3 more than the same number. What is the number? Please show the set up/simplifying/solving of this problem. This is the only one of 25 that has me stumped.
THANKS!!,
Jeff
: I cannot figure out how to set up and solve this word problem. PLEASE HELP!!!!
Here is goes:
"One half of a number is 3 more than the same number. What is the number? Please show the set up/simplifying/solving of this problem. This is the only one of 25 that has me stumped.
THANKS!!,
Jeff

Answer by nerdybill(1042)

One half of a number is 3 more than the same number. What is the number?
.
Let x = "the number"
then
(1/2)x = x+3
multiplying both sides by 2:
x = 2x + 6
subtracting 2x from both sides:
-x = 6
multiplying both sides by -1:
x = -6
Question 146171: I cannot figure out how to set up and solve this word problem. PLEASE HELP!!!!
Here is goes:
"One half of a number is 3 more than the same number. What is the number? Please show the set up/simplifying/solving of this problem. This is the only one of 25 that has me stumped.
THANKS!!,
Jeff
: I cannot figure out how to set up and solve this word problem. PLEASE HELP!!!!
Here is goes:
"One half of a number is 3 more than the same number. What is the number? Please show the set up/simplifying/solving of this problem. This is the only one of 25 that has me stumped.
THANKS!!,
Jeff

Answer by stanbon(18743)

"One half of a number is 3 more than the same number. What is the number?
--------
Let the number be "x".
EQUATION:
(1/2)x = x + 3
(-1/2)x = 3
x = -6
------------
Cheers,
Stan H.